# Tabby: Tabular Data Synthesis with Language Models

### v1

@@ 1 Introduction
From spreadsheets to databases, much of our modern life is encoded in tables. Airplane black boxes, website visitor logs and hospital patient records are just a few examples of this versatile modality. Despite widespread use of tabular data and many calls for improved tabular modeling approaches, this type of data has received less attention in deep learning research than images or text [7, 5, 31].
@@
There have been notable efforts to adapt several model architectures to tabular data, recently focusing on generative adversarial networks (GANs) [33], LLMs [3] and diffusion models [12]. However, because these architectures were each designed with images or text in mind, significant preprocessing must be made to tabular datasets in order to allow their use, likely resulting in lower performance than would be possible for an architecture designed specifically for tabular data.
For these reasons, works including van Breugel and van der Schaar [31] have called for the development of pretrained Large Tabular Models (LTMs) to fill a similar role to text and image foundation models, such as GPT [1] or DALL-E [18]. Unfortunately, the creation of an LTM would require (1) large and diverse tabular pretraining sets which have not yet been curated, (2) a specialized tabular model architecture which has yet to be designed, and (3) a substantial compute resources for pretraining.
This work takes an initial step towards developing an LTM with Tabby, a post-training modification to the standard transformer LLM architecture for enabling tabular data synthesis. After training on text data—but before finetuning on tabular data—Tabby replaces select LLM blocks with Mixture-of-Experts (MoE) layers [27], allowing each data column to be modeled by a dedicated set of parameters in the LLM. The greater expressivity afforded by this change results in higher-fidelity synthetic data. Fine-tuning with our novel Plain technique results in still higher performance. We show that even small Tabby models are capable of outstripping large, non-Tabby LLMs with parameter counts that are orders of magnitude greater.
@@
We introduce our novel tabular training method for LLMs, Plain. Named after its surprising simplicity compared to prior LLM table training approaches, Plain increases data quality in datasets when used together with or independently of Tabby.
@@ 2 Tabby Architecture & Plain Train Method
We now formally introduce our two novel contributions for LLM table synthesis: Tabby is an architecture modification that may be applied to any transformer-based language model (LM) [32], and Plain is a training technique for training any (Tabby or non-Tabby) LM on tabular data. Tabby and Plain are especially powerful together: Tabby increases model expressivity in a way that is specially-suited to tabular data, while Plain allows the model to more easily focus on learning the critical features of a dataset during training.
In Section 2.1, we describe Tabby for tabular or other structured data. In Section 2.2 , we outline the process for training an LM on tabular data using our Plain training technique. Then, in Section 2.3, we provide additional insight into how Tabby models are trained (using Plain, or pre-existing LLM table training techniques such as GReaT [3]) by comparing the training process’s forward pass and loss calculation for a Tabby model with a non-Tabby model.
@@ 2.1 Architecture of Tabby Models
At its core, the Tabby modification is quite simple. The right side of Figure 1 depicts our best-performing Tabby model for tabular data: in this use case, Tabby modifies only the LM’s language modeling head.
Consider a tabular dataset with columns. Let the order of blocks within an arbitrary transformer-based LM be represented as . We apply the MoE technique by replacing an LM block with a vector of blocks. Thus, a Tabby model with one MoE block is represented
The dataset’s -th column is modeled by within .
This technique may be applied to any set of layers within the model. While we focus on the language modeling (LM) head 222Note: Here, “LM head” refers to the language model output layer, distinct from the attention heads in the model’s MLP blocks. in Figure 1 and Section 3 evaluations, we also conduct extensive experiments applying Tabby to the transformer multi-layer perceptrons and attention blocks in Appendix E, visualized in Figure 8. We refer to Tabby models with MoE LM Heads as Tabby Multi-Head (MH) models.
@@ 2.2 The Plain Technique for Fine-Tuning LLMs on Tabular Data
Suppose our training dataset contains rows and the dataset’s column names are denoted by , where the value of the -th row in the -th column is denoted as . To provide the LM with its expected text modality input, we convert the -th row as follows, where <EOS> is the end-of-sequence token and <EOC> is a specialized end-of-column token which we introduce to divide the text between columns:
@@
After converting the tabular dataset in this fashion, an LM is capable of fine-tuning on the dataset in a normal sequence-to-sequence style. Because Plain encodes data in a similar way to prior LLM table training techniques, GReaT and Tabula [3, 38], Plain training does not require more FLOPs than these methods.
@@ 2.3 Tabby Training
Now that we have introduced Tabby and the Plain method to fine-tune LMs on tabular data, we are able to provide further insight into aspects of the training process unique to Tabby. Suppose that we construct a Tabby model from a base LM by replacing one of its blocks with an MoE set . At the beginning of fine-tuning the Tabby model, weights for each block in are initialized to equal the weights of .
The Tabby training process requires only slight modifications compared to other LMs for tabular data. Instead of representing each training row as one string, we convert each row into a length- list of strings as follows:
@@
Internally, the Tabby model begins by training on column with prompt <BOS>, attending to tokens through when predicting the -th token. After computing the loss on column , this column’s tokens are appended to the prompt used to train column . As such, the prompt when training on column is
@@
A favorable side-effect of this training style is that we calculate the losses for each column separately, allowing the performances of each column to be monitored separately and compared, as demonstrated in Section 3.4.
@@ 2.4 Extensions
Synthesis for general structured modalities:
The flexibility in Tabby MoE layer design enables extensions to a variety of structured data, such as hierarchical data. For example, we create a model for nested JSON data by applying Tabby recursively in Figure 7. The JSON structure is preserved inherently in the model: as we show in Section 3.4, this means that Plain’s method of representing data features does not need to be modified to indicate nested features in order to reach synthetic parity with real JSON data.
@@
Because Tabby MoEs contain one block per dataset column, model parameter count is proportionate to the number of data features. In practice, however, techniques such as parameter sharing [23] can drastically reduce the number of parameters required to represent a Tabby model. Additionally, Tabby may be implemented so that only one block in the MOE layer is in memory at a time, resulting in memory requirements identical to a non-Tabby model.
@@ 3 Experimental Results
With our evaluations, we seek to assess the following claims:
Claim 1: Plain-trained Tabby models generate higher-quality data than pre-existing tabular synthesis approaches.
@@
After providing key evaluation setup details in Section 3.0, we compare Tabby to a broad array of prior works on diverse tabular datasets in Section 3.1 to evaluate Claim 1. As Tabby may be applied to any transformer-based LM, we explore Claim 2 for LMs of varying sizes in Section 3.2. To demonstrate Claim 3, we apply Tabby to a nested (JSON) dataset in Section 3.3. Lastly, in Section 3.4, we investigate how Tabby adapts to individual columns within a dataset during finetuning as a demonstration of Claim 4.
@@ 3.0 Setup
Baselines and Comparisons: We include a variety of recent tabular synthesis approaches in our evaluations.
@@
Plain training, our lightweight and simple training paradigm. Plain training is detailed in Section 2.2.
@@
To align with prior works, LLM methods use Distilled-GPT2 as a base model where not otherwise stated.
Non-LLM Approaches: To represent non-LLM tabular synthesis techniques, we include CTGAN [33] and TVAE [33], the leading GAN and VAE approaches, as well as diffusion model Tab-DDPM [12]. Although Tab-DDPM is a SOTA approach to achieving high MLE scores, it does so under strong assumptions—see Figure 4 and Section 4.
@@
| N Train | N Validation | N Test | # Cat. | # Num. | Task | |
| Diabetes [11] | 576 | 57 | 135 | 0 | 8 | C |
| Travel [30] | 715 | 71 | 168 | 4 | 2 | C |
| Adult [2] | 36631 | 3663 | 8548 | 8 | 6 | C |
| Abalone [17] | 3132 | 313 | 732 | 1 | 7 | R |
| Rainfall [34] | 12566 | 1256 | 2933 | 2 | 1 | R |
| House [19] | 15480 | 1548 | 3612 | 0 | 8 | R |
Datasets:
Metrics:
We focus on machine learning efficacy (MLE) [4], the standard metric essential to the quantitative evaluation of synthetic tabular data. In brief, MLE compares the performance of downstream classifiers that were trained using either real or synthetic data.
Our MLE results in the following sections are interpreted as follows: the “Original" downstream classifier, trained using real data, is considered the upper bound and any MLE score higher than the “Original" score is considered the best. If no score surpasses the “Original" score, then any highest score is considered the best.
A visual summary of MLE calculation is provided in Figure 2. In Appendix D, we provide a formal definition of MLE (as well as two more metrics deferred to the Appendices: Discrimination and the Distance to Closest Record).
Aggregation of results: Our evaluation involves a comparison between 9 synthesis methods (including Tabby) across 6 datasets. So, while we do report final scores on each task individually, we would also like to understand which method performs the best across all of the tasks in our evaluation.
To do so, we aggregate MLE scores using performance profile curves [6], which are a robust way to visually compare scores across noisy evaluations in a large number of environments. Performance profiles improve over simpler aggregation techniques, such as averaging scores or computing the average rank of methods across tasks. Specifically, performance profiles are useful when scores for different tasks might be on different scales (which can be an issue with averaging scores), and can take into account methods that are extremely close to the best-performing method on a task without dropping them a full rank (which can be problematic when averaging ranks).
To summarize these curves, we also calculate the area under the performance profile (AUP) scores [25], which serve as a final ranking of methods. In short, the performance of a synthesis method across all six datasets may be represented as just one performance profile curve. Methods with better performance will have higher curves and, therefore, higher AUP scores. As such, the method with highest AUP score is considered the best overall method. Details on performance profiles are in Appendix D.
@@ 3.1 Tabby versus Baseline Synthesis Methods
| Diabetes | Travel | Adult | Abalone | Rainfall | House | |
| Original (Upper Bound) | ||||||
| CTGAN | ||||||
| TVAE | ||||||
| Tab-DDPM | ||||||
| Plain Base DGPT2 | ||||||
| Plain Tabby MH DGPT2 | ||||||
| GReaT Base DGPT2 | N/A* | |||||
| GReaT Tabby MH DGPT2 | * | |||||
| GTT Base DGPT2 | ||||||
| GTT Tabby MH DGPT2 |
@@
Claim 1: Plain-trained Tabby models generate higher-quality data than pre-existing tabular synthesis approaches.
Setup: Table 2 contains MLE results for each dataset. For classification datasets (Diabetes, Travel, Adult), the reported metric is the accuracy of the downstream random forest classifier, while for regression datasets (Abalone, Rainfall, House), we report the coefficient of determination of the downstream random forest regressor.
The “Original" row corresponds to the performance achieved by training the downstream classifier or regressor on real instead of synthetic data. We consider this row to be a performance ceiling for synthetic approaches. Any model and training technique that achieves MLE equal to or better than “Original" is considered to be a top-performing approach and is presented in bold.
Results: We find that Plain-trained Tabby models achieve the highest MLE in datasets. Further, Tabby reaches upper-bound performance on Diabetes, Travel and Adult, indicating that Tabby synthetic data is a capable stand-in for real data in similar scenarios for these datasets.
We also find that Plain is the best-performing technique for training tabular LLMs in almost all cases: for all six datasets, the highest-scoring LLM is trained using Plain. Plain-trained Tabby MH models demonstrate the highest MLE among all LLM architectures and training styles.
@@
Further, we see that the top two synthesis approaches are the two plain-trained models, which surpass the prior SOTA method of Tab-DDPM. Given that these models rely on fewer assumptions than Tab-DDPM, and are simpler to train than the GTT or GReaT LLMs, we find that both Tabby MH and the Plain training technique are powerful advancements for the task of tabular data synthesis.
Comparing Multivariate Modeling Capabilities Across Models:
We further compare the multivariate modeling capabilities of Tab-DDPM, Plain-trained Tabby MH and the prior top-performing LLM-based approach of Great-trained Non-Tabby with TapTap and Tabula in Figure 4. We plot the House dataset’s target column (Median House Value) as a function of its most predictive feature in the dataset (Median Income), for (left to right) real data, plain-trained Tabby MH, non-Tabby GTT and Dab-DDPM.
Tab-DDPM’s plot (leftmost) differs the most from the real data (rightmost) because this model only supports integer-valued regression targets. As a result, both LLM-based approaches more accurately capture the target column’s distribution than Tab-DDPM.
Meanwhile, GReaT sampling (center left) ensures that the target column distribution in the training dataset is replicated in synthetic data. This constraint is enforced by prompting the model with target values selected randomly, based on their frequency in the training data. Accordingly, GReaT models will not generate target values outside those in the training data, which can be undesirable for datasets with few rows or limited target column coverage. In contrast, Plain training (center right) allows the model to generate previously unseen target values. The improved modeling capacity of Tabby over the Non-Tabby model allows Plain’s sampling approach to effectively capture the overall distribution of the target column.

### v2

# Tabby: A Language Model Architecture for Tabular and Structured Data Synthesis

@@ 1 Introduction
From spreadsheets to databases, much of our modern life is encoded in tables. Airplane black boxes, website visitor logs and hospital patient records are just a few examples of this versatile modality. Despite widespread use of tabular data and many calls for improved tabular modeling approaches, this type of data has received less attention in deep learning research than images or text [fang_large_2024, davila_navigating_2024, van_breugel_why_2024].
@@
There have been notable efforts to adapt several model architectures to tabular data, recently focusing on generative adversarial networks (GANs) [xu_modeling_2019], LLMs [borisov_language_2022] and diffusion models [kotelnikov_tabddpm_2022]. However, because these architectures were each designed with images or text in mind, significant preprocessing must be made to tabular datasets in order to allow their use, likely resulting in lower performance than would be possible for an architecture designed specifically for tabular data.
For these reasons, works including van_breugel_why_2024 have called for the development of pretrained Large Tabular Models (LTMs) to fill a similar role to text and image foundation models, such as GPT [achiam2023gpt] or DALL-E [openai_dalle_2021]. Unfortunately, the creation of an LTM would require (1) large and diverse tabular pretraining sets which have not yet been curated, (2) a specialized tabular model architecture which has yet to be designed, and (3) a substantial compute resources for pretraining.
This work takes an initial step towards developing an LTM with Tabby, a post-training modification to the standard transformer LLM architecture for enabling tabular data synthesis. After training on text data—but before finetuning on tabular data—Tabby replaces select LLM blocks with Mixture-of-Experts (MoE) layers [gatedmoe1], allowing each data column to be modeled by a dedicated set of parameters in the LLM. The greater expressivity afforded by this change results in higher-fidelity synthetic data. Fine-tuning with our novel Plain technique results in still higher performance. We show that even small Tabby models are capable of outstripping large, non-Tabby LLMs with parameter counts that are orders of magnitude greater.
@@
We introduce our novel tabular training method for LLMs, Plain. Named after its surprising simplicity compared to prior table training approaches for LLMs, Plain increases data quality in all datasets when used together with or independently of Tabby, compared to the prior SOTA LLM approach.
@@ 2 Tabby Architecture & Plain Train Method
We now formally introduce our two novel contributions for LLM table synthesis: Tabby is an architecture modification that may be applied to any transformer-based language model (LM) [vaswani2017attention], and Plain is a training technique for training any (Tabby or non-Tabby) LM on tabular data. Tabby and Plain are especially powerful together: Tabby increases model expressivity in a way that is specially-suited to tabular data, while Plain allows the model to more effectively fit to the key features of a tabular dataset during training, yielding more realistic data synthesis.
In Section 2.1, we describe Tabby for tabular or other structured data. In Section 2.2 , we outline the process for training an LM on tabular data using our Plain training technique. Then, in Section 2.3, we provide additional insight into how Tabby models are trained (using Plain, or pre-existing LLM table training techniques such as GReaT [borisov_language_2022]) by comparing the training process’s forward pass and loss calculation for a Tabby model with a non-Tabby model.
@@ 2.1 Architecture of Tabby Models
The intuition behind the Tabby modification is simple: we want to allow the model to learn individual columns as distinct–but interdependent–tasks. The right side of Figure 1 depicts our best-performing Tabby model for tabular data, where Tabby modifies only the language modeling head.
To provide a general definition of a Tabby model, consider a tabular dataset with columns. Let the order of blocks within an arbitrary transformer-based LM be represented as . We apply the MoE technique by replacing an LM block with a vector of blocks. Thus, a Tabby model with one MoE block is represented
The dataset’s -th column is modeled by within . In this way, Tabby’s MOE layers use a deterministic, schema-based routing function where the chosen expert is determined by the current column index. The gating function can be expressed mathematically as , where denotes a position within the -th column, and is the -th expert.
This technique may be applied to any set of layers within the model. While we focus on the language modeling (LM) head 222“LM head” refers here to the language model output layer, distinct from attention heads in the MLP blocks. in Figure 1 and Section 3 evaluations, we also conduct experiments applying Tabby to the transformer multi-layer perceptrons and attention blocks in Appendix E. We refer to Tabby models with MoE LM Heads as Tabby Multi-Head (MH) models.
@@ 2.2 The Plain Technique for Fine-Tuning LLMs on Tabular Data
Suppose our trainset contains rows and column names denoted by , such that represents the value of the -th row in the -th column. To provide the LM with its expected text modality input, we convert the -th row as follows, where <EOS> is the end-of-sequence token and <EOC> is a specialized end-of-column token which we introduce to divide the text between columns:
@@
Converting the tabular dataset in this fashion allows an LM to fine-tune on the dataset in a normal sequence-to-sequence style. Because Plain encodes data the same way as prior LLM table training techniques, GReaT and Tabula [borisov_language_2022, zhao_tabula_2023], Plain does not require more FLOPs than prior methods.
@@ 2.3 Tabby Training
Now that we have introduced Tabby and the Plain training method, we are able to provide further insight into aspects of the training process unique to Tabby. Suppose that we construct a Tabby model from a base LM by replacing one of its blocks with an MoE set . At the beginning of fine-tuning the Tabby model, weights for each block in are initialized to equal the weights of . These column experts then diverge from each other over the course of fine-tuning.
The Tabby training process requires only slight modifications compared to other LMs for tabular data. Instead of representing each training row as one string, we convert each row into a list of strings:
@@
Internally, the Tabby model begins by training on column with prompt <BOS>, attending to tokens through when predicting the -th token. After computing the loss on column , this column’s tokens are appended to the prompt used to train column . Backpropagation is performed at the end of a batch of rows. The prompt when training on column is
@@
Because we calculate losses for each column separately, we are able to monitor the performance of each column individually during training. This favorable side-effect is demonstrated in Section 3.4.
@@ 2.4 Extensions
Synthesis for general structured modalities: The flexibility in Tabby MoE layer design enables extensions to a variety of structured datatypes, such as hierarchical data. For example, we create a model for nested JSON data by applying Tabby recursively in Figure 5. The JSON structure is preserved inherently in the model, so that Plain’s method of representing data features does not need to be modified to indicate nested features. As we show in Section 3.4, the combination of Plain and Tabby is the only synthesis approach to reach equal performance to real, non-synthetic data.
@@
Because Tabby MoEs contain one block per dataset column, model parameter count is proportionate to the number of data features. In practice, however, techniques such as parameter sharing [ravanbakhsh2017equivariance] can drastically reduce the number of parameters to represent a Tabby model. Additionally, Tabby may be implemented so that only one block in the MOE layer is in memory at a time, a potential strategy to achieve memory requirements identical to a non-Tabby model.
@@ 3 Experimental Results
Our evaluations seek to assess the following claims:
Claim 1: Plain-trained Tabby models generate higher-quality tabular data than prior approaches.
@@
Claim 5: Tabby’s language modeling capabilities enable it to capture the underlying semantic structure of column values unseen during pretraining, allowing it to generate novel yet realistic values beyond the pretraining distribution.
After providing key evaluation setup details in Section 3.0, we compare Tabby to a broad array of prior works on diverse tabular datasets in Section 3.1 to evaluate Claim 1. As Tabby may be applied to any transformer-based LM, we explore Claim 2 for LMs of varying sizes in Section 3.2. To demonstrate Claim 3, we apply Tabby to a nested (JSON) dataset in Section 3.3. In Section 3.4, we investigate how Tabby adapts to individual columns within a dataset during finetuning as a demonstration of Claim 4. Lastly, we demonstrate Claim 5 in Appendix E.5.
@@ 3.0 Setup
| N Train | N Validation | N Test | # Cat. | # Num. | Task | |
| Diabetes [diabetes] | 576 | 57 | 135 | 0 | 8 | C |
| Travel [tejashvi_tour_2023] | 715 | 71 | 168 | 4 | 2 | C |
| Adult [adult] | 36631 | 3663 | 8548 | 8 | 6 | C |
| Magic [magic_gamma_telescope_159] | 17117 | 1711 | 1902 | 0 | 10 | C |
| Shoppers [online_shoppers_purchasing_intention_dataset_468] | 11097 | 1109 | 1233 | 7 | 10 | C |
| Abalone [abalone] | 3132 | 313 | 732 | 1 | 7 | R |
| Rainfall [zaman_machine_2018] | 12566 | 1256 | 2933 | 2 | 1 | R |
| House [house] | 15480 | 1548 | 3612 | 0 | 8 | R |
Baselines and Comparisons: We evaluate a variety of recent tabular synthesis techniques.
@@
Our lightweight and simple Plain training paradigm, detailed in Section 2.2.
@@
GReaT [borisov_language_2022], which encodes tabular data similarly to Plain, but permutes the orders in which columns are presented in training and imposes some conditional restrictions at sample time. For more details, see Section 4.
@@
GReaT combined with TapTap [zhang_generative_2023] and Tabula [zhao_tabula_2023]. We abbreviate this combination as GTT. TapTap pretrains the LLM on tabular data, while Tabula encodes each categorical column into an ordinal format by replacing each unique column value with an integer.
To align with the prior works [borisov_language_2022, zhang_generative_2023, zhao_tabula_2023], LLM methods use Distilled-GPT2 (DGPT2) [gpt2] as a base model unless otherwise stated.
Non-LLM Approaches: To represent non-LLM tabular synthesis techniques, we include CTGAN+ [zhao_ctab-gan_2022], CTGAN [xu_modeling_2019] and TVAE [xu_modeling_2019], the leading GAN and VAE approaches, as well as diffusion models TabSyn [zhang_mixed-type_2024], TabDiff [shi_tabdiff_2025], Tab-DDPM [kotelnikov_tabddpm_2022] and Forest Diffusion [jolicoeur-martineau_generating_2024]. Although diffusion models are a SOTA approach to achieving high MLE scores, they do so under strong assumptions and are incompatible with many tabular datasets—see Figure 4 and Section 4.
@@
Datasets: We evaluate Tabby on eight common tabular datasets, which are summarized in Table 1. The majority of these datasets are standard for the evaluation of tabular synthesis techniques, allowing for easy comparison with prior approaches. For more information on these datasets, see Appendix B.
Metrics: We focus on machine learning efficacy (MLE) [mle], the standard metric for quantitative evaluation of synthetic tabular data. In brief, MLE compares the performance of downstream classifiers that were trained using either real or synthetic data.
Our MLE results in the following sections are interpreted as follows: the downstream classifier that is trained using non-synthetic, real data is considered the upper bound and any MLE score higher than this “Non-Synthetic” classifier’s score is considered the best. If no score surpasses the “Non-Synthetic” score, then any highest score is considered the best. Figure 2 summarizes the MLE calculation process. In Appendix E, we provide a formal definition of MLE, as well as several more metrics including Shape, Trend [shi_tabdiff_2025], Distance to Closest Record (DCR) and Membership Inference Attack [yao2025dcrdelusionmeasuringprivacy] performance. These metrics additionally compare the synthetics’ abilities to preserve trainset distributions without memorizing the trainset.
Aggregation of results: Our evaluation involves a comparison between 14 synthesis methods (including Tabby) across eight datasets. So, while we do report final scores on each task individually, we would also like to understand which method performs the best across all of the tasks in our evaluation.
To do so, we aggregate MLE scores using performance profile curves [dolanBenchmarkingOptimizationSoftware2002], a robust way to visually compare scores across noisy evaluations in a large number of environments. Performance profiles improve over simpler aggregation techniques, such as averaging scores or computing the average rank of methods across tasks. Specifically, performance profiles are useful when scores for different tasks might be on different scales (which can be an issue with averaging scores), and can take into account methods that are extremely close to the best-performing method on a task without dropping them a full rank (which can be problematic when averaging ranks).
To summarize these curves, we also calculate the area under the performance profile (AUP) scores [automldecathlon], which serve as a final ranking of methods. In short, the performance of a synthesis method across all eight datasets may be represented as just one performance profile curve. Methods with better performance will have higher curves and, therefore, higher AUP scores. As such, the method with highest AUP score is considered the best overall method. Details on performance profiles are in Appendix D.
@@ 3.1 Tabby versus Baseline Synthesis Methods
| Diabetes | Travel | Adult | Magic | Shoppers | Abalone | Rainfall | House | Best | |
| Non-Synthetic | |||||||||
| CTGAN | 0 | ||||||||
| CTGAN+ | 0 | ||||||||
| TVAE | 0 | ||||||||
| CLLM | N/A* | 2 | |||||||
| TabSyn | 1 | ||||||||
| TabDiff | 4 | ||||||||
| Forest Diffusion | 1 | ||||||||
| Tab-DDPM | 3 | ||||||||
| Plain Base | 4 | ||||||||
| Plain Tabby MH | 5 | ||||||||
| GReaT Base | N/A* | 0 | |||||||
| GReaT Tabby MH | * | 1 | |||||||
| GTT Base DGPT2 | 1 | ||||||||
| GTT Tabby MH | 1 |
@@
Claim 1: Plain-trained Tabby models generate higher-quality tabular data than prior approaches.
Setup: Table 2 lists MLE for each dataset. For classification datasets (Diabetes, Travel, Adult), the reported metric is the accuracy of the downstream random forest classifier, while for regression datasets (Abalone, Rainfall, House), we report the coefficient of determination of the downstream random forest regressor.
The “Non-Synthetic" row corresponds to the performance achieved by training the downstream classifier or regressor on real instead of synthetic data. We consider this row to be a performance ceiling for synthetic approaches. Any model and training technique that achieves MLE equal to or better than “Non-Synthetic" is considered to be a top-performing approach and is presented in bold.
Results: We find that Plain-trained Tabby models achieve the highest MLE in datasets. Further, Tabby reaches upper-bound performance on Diabetes, Travel, Adult, Magic and Shoppers, indicating that Tabby synthetic data is a capable stand-in for real data in similar scenarios for these datasets.
We also find that Plain is the best-performing technique for training tabular LLMs in almost all cases: for all eight datasets, the highest-scoring LLM is trained using Plain. Plain-trained Tabby MH models demonstrate the highest MLE among all LLM architectures and training styles.
@@
Further, we see that the top two synthesis approaches are the two Plain-trained models, which surpass the prior SOTA method of TabDiff. Given that these models rely on fewer assumptions than diffusion approaches, and are simpler to train than the GTT or GReaT LLMs, we find that both Tabby MH and Plain training are powerful advancements for the task of tabular data synthesis.
Comparing Multivariate Modeling Capabilities:
We further compare the multivariate modeling capabilities of Tab-DDPM, Plain-trained Tabby MH and the prior top-performing LLM-based approach of GReaT-trained Non-Tabby with TapTap and Tabula in Figure 4. We plot the House dataset’s target column (Median House Value) as a function of its most predictive feature in the dataset (Median Income), for (left to right) real data, Plain Tabby MH, non-Tabby GTT and Tab-DDPM.
Tab-DDPM’s plot (left) differs the most from the real data (right) because the Tab-DDPM model only supports integer-valued regression targets. Accordingly, both LLM-based approaches more accurately capture the target column’s distribution than Tab-DDPM.
Meanwhile, GReaT sampling (center left) constrains that the target column distribution in the training dataset is replicated in synthetic data, by prompting the model with target values selected randomly based on their frequency in the training data. Accordingly, GReaT models will not generate target values outside those in the training data, which can be undesirable for datasets with few rows or limited target column coverage. In contrast, Plain training (center right) allows the model to generate previously unseen target values. The improved modeling capacity of Tabby over the Non-Tabby model allows Plain’s sampling approach to effectively capture the overall target column distribution, demonstrating the advantages of using both Tabby and Plain together.